IB] 0.0 0.2 0.4 0.6 0.8 What is the value of the equilibrium constant for this IAI 1.0 0.8 0.6 0.4 0.2 reaction? (0.50 pt) (a) 0.25 (b) 0.66 c) 1.0 (d) 1.5 (e) 4.0 8. Consider the process: N20(g, 1 atm) N20g, latm) O(g, 1 atm) at 25.0°C? (0.5 pt) (a) What is AS (Use the necessary thermodynamic data from the table to solve this problem) (b) What is AS (Use the necessary thermodynamic data from the tables to solve this problem) (0 Nse surr (c) Use your results from parts (a) and (b) to determine whether this process is spontaneous at this temperature and standard conditions? Explain briefly. (0.5 pt) solubility questions at 298 K. What is AGO at 298 K for the precipitation of Al(oH)3 a to the equation below: (1.5 pts)
media%2F1d5%2F1d5e94c1-6289-49ef-938a-d4
I need help on 8(a)
Answer #1

8. (a) ΔSosys = ΣnSo (Products) - ΣnSo (Reactants)

where n = No. of moles, So = Standard entropy

From the table, So of N2O = 219.7 J/(mol K), So of N2 = 191.5 J/(mol K), So of O = 160.9 J/(mol K)

From the reaction, n of N2O = 1 mol, n of N2 = 1 mol, n of O = 1 mol

ΣnSo (Products) = nSo (N2) + nSo (O) = [1 mol X 191.5 J/(mol K)] + [1 mol X 160.9 J/(mol K)] = 352.4 J/K

ΣnSo (Reactants) = nSo (N2O)= [1 mol X 219.7 J/(mol K)] = 219.7 J/K

ΔSosys = ΣnSo (Products) - ΣnSo (Reactants) = 352.4 J/K - 219.7 J/K = 132.7 J/K